OpenStudy (anonymous):

how you Evaluate this: −6 − (−2)

OpenStudy (apoorvk):

When two adjacent 'minus' signs are present, they turn into a 'plus' sign. So, -6 -(-2) = -6 + 2 = ?? what would this be?

OpenStudy (anonymous):

-6 + 2 -4 Two negatives make a positive :)

OpenStudy (anonymous):

-8?

OpenStudy (anonymous):